# TinyVLA: Towards Fast, Data-Efficient Vision-Language-Action Models for Robotic Manipulation

### v4

@@ Abstract
Vision-Language-Action (VLA) models have shown remarkable potential in visuomotor control and instruction comprehension through end-to-end learning processes. However, current VLA models face significant challenges: they are slow during inference and require extensive pre-training on large amounts of robotic data, making real-world deployment difficult. In this paper, we introduce a new family of compact vision-language-action models, called TinyVLA, which offers two key advantages over existing VLA models: (1) faster inference speeds, and (2) improved data efficiency, eliminating the need for pre-training stage. Our framework incorporates two essential components to build TinyVLA: (1) initializing the policy backbone with robust, high-speed multimodal models, and (2) integrating a diffusion policy decoder during fine-tuning to enable precise robot actions. We conducted extensive evaluations of TinyVLA in both simulation and on real robots, demonstrating that our approach significantly outperforms the state-of-the-art VLA model, OpenVLA, in terms of speed and data efficiency, while delivering comparable or superior performance. Additionally, TinyVLA exhibits strong generalization capabilities across various dimensions, including language instructions, novel objects, unseen positions, changes in object appearance, background variations, and environmental shifts, often matching or exceeding the performance of OpenVLA. We believe that TinyVLA offers an interesting perspective on utilizing pre-trained multimodal models for policy learning. Our project is at https://tiny-vla.github.io.
@@ I Introduction
Training multitasking robot imitators to operate in complex and uncertain environments faces considerable challenges due to limited data and the difficulty of learning physical motion [1, 2, 3, 4]. Moreover, traditional robot models struggle to adapt to new scenes and tasks and are easily affected by distractors, lighting conditions, and background changes [5, 6]. Modern methods typically leverage off-the-shelf Large Language Models (LLMs) [7, 8, 9, 10, 11, 12] for scene descriptions to generate object affordance, location, or heatmaps, followed by a predefined motion planner to complete the tasks [13, 14, 15, 16, 17, 18].
@@
In both simulations and real-world settings, our method demonstrates superior performance in multi-task learning compared to the baseline. For instance, in real-world experiments, TinyVLA-H achieves a 25.7% higher success rate than OpenVLA, while using 5.5 times fewer parameters. In bimanual real-robot experiments, we find that OpenVLA, which heavily relies on OpenX robot data pretraining, struggles to perform in bimanual settings due to OpenX consisting only of single-arm data. In contrast, TinyVLA-H significantly outperforms OpenVLA in these tasks. Additionally, we observed that TinyVLA generalizes well across diverse settings, including observational and spatial generalization, often matching or even surpassing OpenVLA in certain cases. We believe that TinyVLA offers a novel perspective to building vision-language-action models for embodied control.
@@ II Related Works
Multimodal Models connect vision and language and extend the reasoning ability of LLMs to process with multimodal input. Numerous works have been proposed in this direction [22, 23, 24, 25, 26, 27, 28, 29, 30]. These MLLMs typically have parameters ranging from 7B to 70B, making the inference cost-prohibitive and limiting the accessibility of MLLMs to a wider audience. Recently, a select number of studies [31, 32, 33, 34, 35, 36, 37]have delved into the exploration of efficient multimodal, with a number of parameters less than 3B, from diverse angles. Robot learning [38, 39, 19, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] is an crucial topic in the robotics. A number of works introduce vision-language models to the domain of embodied control, these representative works include RT-2 [19] and OpenVLA [20]. Our paper proposes a compact version of VLA models that does not require large-scale robot data pretraining and enables fast inference.
@@ III Method
This section gives a comprehensive overview of our proposed TinyVLA. TinyVLA encompasses several crucial designs: 1) We adopt a pre-trained Multimodal model as the initialization of a policy network; 2) During training the robot data, we freeze the pre-trained parts and utilize the parameter-efficient fine-tuning technique LoRA [62], where the trainable parameters account for only 5% of the entire model; 3) We introduce a policy decoder that concatenated to pre-trained multimodal model through a simple but efficient linear projection and output the executable action of the robot. An illustration of TinyVLA is given in Figure 2.
@@ III-A Building TinyVLA with Efficient Vision-Language Models
Figure 2: Model architecture.The left image illustrates the VLM pretraining pipeline, whereas the right image demonstrates the process of training TinyVLA using robotic data.
The initial step involves acquiring pre-trained multimodal language models. While existing works typically focus on vision-language models with over three billion parameters, we trained a more compact vision-language model with parameters ranging from 40 million to 1.4 billion. Our model utilizes Pythia [63] as the language model backend. We then followed the training pipeline of LLaVA [25], using their vision-language dataset to train this family of VLMs. For robot data fine-tuning, we retained all modules from our VLM, including the visual backbone and the vision-language alignment module.
@@ III-B Robot Data Finetuning for Manipulation
Learning action with diffusion policy decoder. We need a way to represent the action space to control the robot. One method is to use discrete tokenization for the actions, as has been done in RT-2. However, using tokenization for continuous or high-dimensional data has proven to be extremely challenging for training [64], requires a huge amount of data [65, 66], and tends to converge to a single state [67]. Therefore, instead of converting actions into token space, we leverage a policy head to further learn the action space of the robot.
Specifically, we leverage the Diffusion Policy (DP) [5]. DP formulates robot policies using Denoising Diffusion Probabilistic Models (DDPMs) [68] which mainly involves two processes: adding noise and denoising. During training, a random value $K$ is selected from 0 to $N$, indicating the number of times Gaussian noise is added to the original actions. Subsequently, DP denoises the noisy actions by predicting the added noise and minus it to acquire denoised actions. In the inference phase, starting from a pure noise $a^{N}$ sampled from Gaussian distribution, DP generates a series of intermediate actions, $\{a^{n},a^{n-1},...,a^{1}\}$, until a desired noise-free output $a^{0}$ is formed. The whole model structure is shown in Figure 2 right, we directly connect DP and multimodal model backbone through two simple Linear projections and one LayerNorm. The multimodal model backbone jointly encodes the current observations and language instructions to generate a multimodal embedding that controls the denoising process of DP as conditions.
@@ IV-A2 Real Robot Setup
Data collection. We collect the dataset through teleoperation. We record the RGB stream from two camera views and robot states e.g., joint position during the whole robot control process. TinyVLA predicts the 6D pose, including position $(x,y,z)$ and rotation $(roll,pitch,yaw)$. For all the tasks we do not add additional distractors except in the remove the lid of the box task, in order to better evaluate the model’s generalization capability to distractors. In total, we collected 100 trajectories for each task to balance data distribution across all 5 tasks.
@@
Figure 5: View Generalization. We evaluated the view generalization capability of our model in a new environment, which we designed to be as consistent with the training environment as possible. We tested 3 tasks respectively, and the results under 8 viewing angle changes (the two cameras each correspond to 4 changes).
@@ IV-B Experimental Results on Multi-Task Learning
Real-world experimental results. The experimental results are shown in Table II. We evaluate each model 20 trials per task in single arm setting. We report the mean and standard deviation of success rates across 3 checkpoints. Notably, TinyVLA-H attained a 98.3% success rate in flipping a mug, stacking cubes, and a 90% success rate in place tennis, leading a large margin over other baselines. It is also worth mentioning that both place tennis and stack cubes are position-sensitive tasks. Also, both objects are rigid bodies, which makes the task challenging. Specifically, regarding the average success rate over five tasks, the result of TinyVLA-H surpasses OpenVLA by 25.7%.
@@ IV-E Experiments on Generalization
Figure 6: Background Generalization. We utilized six different backgrounds, testing three of them on Task a and the remaining three on Task b.
Figure 7: Distractor & Illumination Generalization. In the distractor settings, Level L1 involves the addition of irrelevant objects such as books and cups, which are unrelated to the task. Level L2 involves the inclusion of identical cubes in various colors, adding complexity to the visual environment. Regarding illumination settings, Level L1 represents conditions with reduced lighting, while Level L2 describes scenarios with minimal lighting.
@@ IV-G Visual Generalization
Figure 8: Object & Appearance generalization. For object generalization, we replace the objects with previously unseen ones that have different shapes or colors. For appearance generalization, we only alter the colors of the objects.
Figure 9: Spatial generalization. We conducted evaluations at multiple positions thoroughly outside the training zone on two position-sensitive tasks:place tennis and flip mug.
V Conclusion
@@
VI Acknowledgements
We sincerely thank Yanjie Ze for contributions in discussions and paper review. Our also thank Yong Wang for his assistance with the hardware setup.

### v5

@@ Abstract
Vision-Language-Action (VLA) models have shown remarkable potential in visuomotor control and instruction comprehension through end-to-end learning processes. However, current VLA models face significant challenges: they are slow during inference and require extensive pre-training on large amounts of robotic data, making real-world deployment difficult. In this paper, we introduce a new family of compact vision-language-action models, called TinyVLA, which offers two key advantages over existing VLA models: (1) faster inference speeds, and (2) improved data efficiency, eliminating the need for pre-training stage. Our framework incorporates two essential components to build TinyVLA: (1) initializing the policy backbone with robust, high-speed multimodal models, and (2) integrating a diffusion policy decoder during fine-tuning to enable precise robot actions. We conducted extensive evaluations of TinyVLA in both simulation and on real robots, demonstrating that our approach significantly outperforms the state-of-the-art VLA model, OpenVLA, in terms of speed and data efficiency, while delivering comparable or superior performance. Additionally, TinyVLA exhibits strong generalization capabilities across various dimensions, including language instructions, novel objects, unseen positions, changes in object appearance, background variations, and environmental shifts, often matching or exceeding the performance of OpenVLA. We believe that TinyVLA offers an interesting perspective on utilizing pre-trained multimodal models for policy learning.
Index Terms:
AI-based method, Deep Learning in Grasping and Manipulation.
@@ I Introduction
Training multitasking robot imitators to operate in complex and uncertain environments faces considerable challenges due to limited data and the difficulty of learning physical motion [1, 2]. Moreover, traditional robot models struggle to adapt to new scenes and tasks and are easily affected by distractors, lighting conditions, and background changes [3, 4]. Modern methods typically leverage off-the-shelf Large Language Models (LLMs) [5, 6] for scene descriptions to generate object affordance, location, or heatmaps, followed by a predefined motion planner to complete the tasks [7, 8].
@@
In both simulations and real-world settings, our method demonstrates superior performance in multi-task learning compared to the baseline. For instance, in real-world experiments, TinyVLA-H achieves a 25.7% higher success rate than OpenVLA, while using 5.5 times fewer parameters. In bimanual real-robot experiments, we find that OpenVLA, which heavily relies on OpenX robot data pretraining, struggles to perform in bimanual settings due to OpenX consisting only of single-arm data. In contrast, TinyVLA-H significantly outperforms OpenVLA in these tasks. Additionally, we observed that TinyVLA generalizes well across diverse settings, including observational and spatial generalization, often matching or even surpassing OpenVLA in certain cases.
Our contribution are the three folds:
  • We introduce a novel VLA architecture that combines lightweight vision-language models with a diffusion model, enabling fast inference, strong performance, and excellent generalization capabilities.
  • We conducted extensive experiments in both simulated and real-world settings, encompassing single-arm and bimanual robot setups, to validate the effectiveness of our method.
  • We demonstrate that strong VLA models can be trained without requiring large-scale robotic datasets, achieving both data-efficiency and high performance.
We believe that TinyVLA offers a novel perspective to building vision-language-action models for embodied control.
@@ II Related Works
Vision-language models (VLMs). VLMs connect vision and language and extend the reasoning ability of LLMs to process with multimodal input. Numerous works have been proposed in this direction [12, 13, 14, 15]. These MLLMs typically have parameters ranging from 7B to 70B, making the inference cost-prohibitive and limiting the accessibility of MLLMs to a wider audience. Recently, a select number of studies [16, 17] have delved into the exploration of efficient multimodal, with a number of parameters less than 3B, from diverse angles. These models run efficiently,
Vision-language models for robot learning. Robot learning [9, 18, 19, 20, 21] is an crucial topic in the robotics. A number of works introduce vision-language models to the domain of robot learning, including using VLMs for high-level planning [22], task decomposition [7], and formulate VLMs as a robot action predictor with end-to-end training [9, 10, 18]. In this work, we explore two perspectives on using VLM as a robot action predictor, 1) how to use a more lightweight and fast VLM and 2) how to replace the autoregression model with a diffusion model.
Multi-task robot learning. Recent advances in multi-task robotic manipulation have yielded significant progress in executing complex tasks and generalizing to novel scenarios [23, 24, 25]. Leading methods often leverage extensive interaction data  [26] to train multi-task models. For example, RT-1 [23] underscores the benefits of task-agnostic training and RT-2 [9] trains with mixed robot data and image-text pairs. PerAct [27] encodes language goals and shows its effectiveness in real robot experiments. Octo [28] uses cross-embodiment data for pertaining. This paper proposes a new approach to learning multi-task policy using a new form of vision-language-action models.
@@ III Method
This section gives a comprehensive overview of our proposed TinyVLA. TinyVLA encompasses several crucial designs: 1) We adopt a pre-trained VLM as the initialization of a policy network; 2) During training the robot data, we freeze the pre-trained parts and utilize the parameter-efficient fine-tuning technique LoRA [29], where the trainable parameters account for only 5% of the entire model; 3) We introduce a policy decoder that concatenated to pre-trained multimodal model through a simple but efficient linear projection and output the executable action of the robot. An illustration of TinyVLA is given in Figure 2.
@@ III-A Building TinyVLA with Efficient Vision-Language Models
Figure 2: Model architecture. The left image illustrates the VLM pretraining pipeline, whereas the right image demonstrates the process of training TinyVLA using robotic data. We adopt diffusion policy as our policy head.
The initial step involves acquiring pre-trained vision-language models (VLM). While existing works typically focus on vision-language models with over three billion parameters, we trained a more compact vision-language model with parameters ranging from 70 million to 1.4 billion. Our model utilizes Pythia [30] as the language model backend. We then followed the training pipeline of LLaVA [13], using their vision-language dataset to train this family of VLMs. For robot data fine-tuning, we retained all modules from our VLM, including the visual backbone and the vision-language alignment module.
@@ III-B Robot Data Finetuning for Manipulation
Learning action with diffusion policy decoder. We need a way to represent the action space to control the robot. One method is to use discrete tokenization for the actions, as has been done in RT-2. However, using tokenization for continuous or high-dimensional data has proven to be extremely challenging for training [31], requires a huge amount of data [32, 33], and tends to converge to a single state [34]. Therefore, instead of converting actions into token space, we leverage the Diffusion Policy(DP) [3] as our policy head. DP formulates robot policies using Denoising Diffusion Probabilistic Models (DDPMs) [35] which predicts the noise instead of direct actions.
The whole framwork is illustrated in Figure 2(right). And the pipeline can be splited into 3 steps. First, the visual-language model (VLM) backbone encodes raw observations and language instructions into multimodal embedding vectors. To handle the inherent variability in input sequence lengths, we employ an adaptive pooling layer followed by layer normalization, producing fixed and compact feature representations. Then, these normalized features are subsequently concatenated with the robot’s proprioceptive state vector. The combined representation is processed through a 3-layer multilayer perceptron (MLP), generating conditional embeddings for the standard training process of DP. Finally, we utilize the standard training method of DP to train the whole VLA model. To preserve the intrinsic knowledge of the pretrained VLM, we implement different training strategies: The VLM is fine-tuned using low-rank adaptation(LoRA), while the DP head undergoes full-parameter training.
@@ IV-A2 Real Robot Setup
Data collection. We collect the dataset through teleoperation. We record the RGB stream from two camera views and robot states e.g., joint position during the whole robot control process. We record the robot gripper’s width as a value between 0 and 1, where 0 represents fully closed and 1 represents fully open. TinyVLA predicts 7-dimensional actions, including position $(x,y,z)$, rotation $(roll,pitch,yaw)$ and $(gripper\_width)$. For all the tasks we do not add additional distractors except in the remove the lid of the box task, in order to better evaluate the model’s generalization capability to distractors. In total, we collected 100 trajectories for each task to balance data distribution across all 5 tasks.
@@
Figure 5: View Generalization. We evaluated the view generalization capability of our model in a new environment, which we designed to be as consistent with the training environment as possible. We tested 3 tasks respectively, and the results under 8 viewing angle changes (the two cameras each correspond to 4 changes).For each specified camera view(e.g. Left Camera -30 degrees), we evaluate each model for 2 trials.
@@ IV-B Experimental Results on Multi-Task Learning
Real-world experimental results. The experimental results are shown in Table II. We evaluate each model 20 trials per task in single arm setting. We report the mean and standard deviation of success rates across 3 checkpoints. Notably, TinyVLA-H attained a 98.3% success rate in flipping a mug, stacking cubes, and a 90% success rate in place tennis, leading a large margin over other baselines. Besides, the performance increases drastically from TinyVLA-S to TinyVLA-H which is adhere to scaling law. What’s more, regarding the average success rate over five tasks, the result of TinyVLA-H surpasses OpenVLA by 25.7%.
@@ IV-E Experiments on Generalization
Figure 6: Background Generalization. We utilized six different backgrounds, testing three of them on Task a and the remaining three on Task b. For each background, we evaluate each model for two trials.
Figure 7: Distractor & Illumination Generalization. For distractor settings, Level L1 involves the addition of objects such as books and cups, which are unrelated to the task. Level L2 involves the inclusion of identical cubes in various colors, adding complexity to the visual environment. For illumination settings, Level L1 represents conditions with reduced lighting, while Level L2 describes scenarios with minimal lighting. For each specified setting (i.e., Distractor L1), we evaluate each model for six trials.
@@
Figure 8: Object & Appearance generalization. For object generalization, we replace the objects with previously unseen ones that have different shapes or colors. For appearance generalization, we only alter the colors of the objects. For each specified setting, we evaluate each model for 6 trials.
Figure 9: Spatial generalization. We conducted evaluations at multiple positions thoroughly outside the training zone on two position-sensitive tasks:place tennis and flip mug. For each out-of-distribution positions, we evaluate each model for 2 trials.
@@ IV-G Visual Generalization
Figure 10: Types of failure for TinyVLA with different sizes of pre-trained vision-language models.
V Ablation Study
V-A Trade-off between size of VLM and TinyVLA’s Performance
Our main experiments (Table II and Table III) demonstrate that our method adheres to the scaling law: as model size increases, the average success rate across tasks improves accordingly. To understand the underlying reasons for this trade-off between model size and performance, we conducted a failure case analysis. We evaluated three TinyVLA variants: TinyVLA-0.4B, TinyVLA-1.3B, and TinyVLA-3B. The first two were used in the main experiments, while TinyVLA-3B utilizes the pre-trained PaliGemma model [45]. Each model was tested on one of four tasks: PlaceTennis, FlipMug, StackCubes on a Franka robot, and PlaceTennisBag on a bimanual robot. Each task was evaluated six times, and the total number of failures was recorded. The results are presented in Figure 10).
Our analysis reveals that VLM size significantly impacts task success. For example, TinyVLA-0.4B failed three times due to misinterpreting instructions, likely because the smaller VLM has limited language comprehension capabilities. This issue was resolved when we increased the model size to 1.3B. Furthermore, increasing the model size mitigated failures related to inaccurate positioning and reaching incorrect target locations. This improvement can be attributed to the use of models like PaliGemma, which are trained on localization data and possess richer visual feature representations, leading to enhanced localization abilities.
V-B Choice of Policy Model
Our TinyVLA model exhibits strong performance and generalization capabilities. This success is largely attributed to the integration of a pre-trained Vision-Language Model (VLM) with a diffusion model. However, this raises a crucial question: how essential is the diffusion model to this architecture? Could alternative methods achieve comparable results? To investigate this, we compared TinyVLA’s performance with two different policy networks: a vanilla multi-layer perceptron (MLP) commonly used for behavior cloning, and an action chunking transformer (ACT) [46] known for generating stable and smooth actions, as shown in Table V. Our findings demonstrate that the diffusion model significantly outperforms both ACT and MLP. Notably, the MLP-based approach failed entirely across all tasks, likely due to the limited capacity of the MLP layers compared to the VLM, hindering effective optimization. While ACT demonstrated some success, our diffusion model achieved a considerably higher average success rate across all tasks. These results underscore the significant advantage of employing a diffusion model over alternative policy networks in our TinyVLA framework.
V-C Which Part of TinyVLA makes it Fast?
A key advantage of TinyVLA is its lightweight design and superior speed compared to OpenVLA. As demonstrated in Figure 1, our largest model, TinyVLA-H, achieves a higher average success rate than OpenVLA while utilizing 5.5 times fewer parameters and operating 20 times faster. This section analyzes the primary contributor to this significant speed advantage. Specifically, we replaced the Prismatic-7B VLM backbone in OpenVLA with the same architecture in TinyVLA. We demonstrate the comparison in Table IV. We observed a reduction in per-action prediction time from 292ms to 140ms. Despite this 2x speed increase, OpenVLA remains 10 times slower than TinyVLA-H, with a similar number of parameters. This result highlights that the speed of TinyVLA stems not only from utilizing a smaller VLM but also from employing a diffusion model for action prediction. This approach avoids the computationally expensive autoregressive generation of action tokens, improving test-time speed.
TABLE IV: Inference latency is measured quantitatively and reported in milliseconds (ms). The reported values represent the time required for a single action prediction by the OpenVLA-1B/7B and TinyVLA-1B models. Experiments are conducted using a single A6000 GPU.
Inference Latency on A6000 GPU
OpenVLA-7B $\rightarrow$ OpenVLA-1B TinyVLA-1B
292 $ms$rightarrow 140 $ms$ 14 $ms$
TABLE V: Ablation study on different choices of policy model. We choose TinyVLA-H as our base model and replace the diffusion model with ACT model [46] and vanilla MLP head. We report the success rate on 5 real robot tasks.
Policy Head PlaceTennis FlipMug StackCubes CloseDrawer OpenBox
Multi-Layer Perceptron 0 $\pm$ 0 0 $\pm$ 0 0 $\pm$ 0 0 $\pm$ 0 0 $\pm$ 0
Action Chunking Transformer 13.3$\pm$0.1 8.3$\pm$0.1 8.3$\pm$0.3 13.3$\pm$0.1 23.3$\pm$0.1
Diffusion Model 90$\pm$0.2 98.3$\pm$0.1 98.3$\pm$0.1 96.7$\pm$0.3 86.7$\pm$0.1
VI Conclusion
@@
Acknowledgement
We sincerely thank Yanjie Ze for contributions in discussions and paper review. We also thank Yong Wang for his assistance with the hardware setup.